OpenStudy (anonymous):

Simplify -16+10-14-(-17) I got -9

OpenStudy (anonymous):

-16+10=-6 -6+-14=-20 -20+17= -3 So I got -3

OpenStudy (anonymous):

Ty

OpenStudy (anonymous):

Your Welcome :)